OpenStudy (tylerd):

find unit vectors that give steepest ascent and descent at point, find vector that points in direction of no change.

OpenStudy (tylerd):

\[f(x,y)=x^4-x^2y+3y^2+7\] point (-1,1)

OpenStudy (tylerd):

\[gradient = <4x^3-2xy , x^2+6y>\]

OpenStudy (jango_in_dtown):

no wrong

OpenStudy (jango_in_dtown):

<4x^3-2xy,-x^2+6y>

OpenStudy (tylerd):

alright, typo, next step?

OpenStudy (jango_in_dtown):

next plug in x=-1 and y=1

OpenStudy (jango_in_dtown):

you get <-2,5>

OpenStudy (tylerd):

yep

OpenStudy (jango_in_dtown):

then find the unit vector by just diving by modulus and you get???

OpenStudy (tylerd):

so for steepest ascent it goes in the same direction as the gradient?

OpenStudy (jango_in_dtown):

\[<-2/\sqrt{29},5/\sqrt{29}>\]

OpenStudy (jango_in_dtown):

right and for the steepest descent just multiply the ascent by -1 and you get??

OpenStudy (tylerd):

steepest decent i just times each component by -1?

OpenStudy (tylerd):

ya

OpenStudy (tylerd):

now for 0 change?

OpenStudy (jango_in_dtown):

and last part you know? you need to find a vector perpendicular to the gradient

OpenStudy (tylerd):

so, dot product = 0?

OpenStudy (jango_in_dtown):

Let the vector by <a,b> then a(-2/sqrt{29})+b(5/sqrt{29})=0 this gives -2a+5b=0 gives a=(5/2)b

OpenStudy (jango_in_dtown):

in particular take b=2 then a=5 therefore (5,2) is such a vector now you may divide by the modulus and get <5/sqrt(29),2/sqrt(29)> the required vector

OpenStudy (jango_in_dtown):

@TylerD

OpenStudy (tylerd):

hmm, for some reason the answer was -5,-2

OpenStudy (jango_in_dtown):

yeah it is also a choice..

OpenStudy (tylerd):

ya i think i c

OpenStudy (jango_in_dtown):

a=(5/2)b if you put b=-2 then a becomes -5

OpenStudy (jango_in_dtown):

you didnt give me options

OpenStudy (jango_in_dtown):

actually there exists infinite number of solutions

OpenStudy (jango_in_dtown):

since a direction is given by a line and a line is composed of infinite number of points so the solution is not unique . the solution is given by the equation a=(5/2)b where you will get value of a if you plug in b

OpenStudy (jango_in_dtown):

@TylerD